OpenStudy (anonymous):

3 4 over 7 times 2 4 over 5

OpenStudy (jack1):

34 or 3x4?

OpenStudy (anonymous):

its a mixed number

OpenStudy (jack1):

...what does that mean, sorry?

geerky42 (geerky42):

He means \(3 \dfrac{4}{7} \times 2 \dfrac{4}{5}\)

OpenStudy (anonymous):

Thank You! i couldnt figure out how to express it

OpenStudy (jack1):

\[\frac{ 25 }{ 7 } \times\frac{ 14 }{ 5 }??\]

OpenStudy (jack1):

times the tops across and then the bottoms across

OpenStudy (jack1):

25 x 14 = over 7 x 5 =

OpenStudy (jack1):

so what did u get as your final answer?

OpenStudy (anonymous):

350 over 35

OpenStudy (anonymous):

10

OpenStudy (jack1):

perfect

OpenStudy (anonymous):

thank you!!

OpenStudy (jack1):

welcome